a 50% juice is to be mixed witha 30% juice to get 200 liters of a 45% juice mixture how many liters of each must be mixed
Answer by stanbon(75887)

a 50% juice is to be mixed witha 30% juice to get 200 liters of a 45% juice mixture how many liters of each must be mixed
----
Equation:
active + active = active
0.50x + 0.30(200-x) = 0.45*200
---------------
50x + 30*200 - 30x = 45*200
20x = 15*200
x = (3/4)200
x = 150 liters (amt. of 50% juice needed)
200-x = 50 liters (amt. of 30% joice needed)
---------
Cheers,
Stan H.
-----------